OpenStudy (anonymous):

What is the value of x in the equation 2(2x − 5 − 4) = 160 − 17? 40.25 62.25 80.25 143.0

OpenStudy (anonymous):

Start by combining like terms, like the -5 and -4, and the 160 and -17

OpenStudy (anonymous):

ok 2(2x-9)=153

OpenStudy (anonymous):

Okay, now divide both sides by 2

OpenStudy (anonymous):

2x-9=72.5?

OpenStudy (anonymous):

you mean distribute?

OpenStudy (anonymous):

Oh wait, you made a mistake when you subtracted the 17, it should have been 143

OpenStudy (anonymous):

Then add nine to both sides, and divide by two, and you should have your answer

OpenStudy (anonymous):

got confused.. ops

OpenStudy (anonymous):

So back to the beginning. 2(2x-5-4)=160-17

OpenStudy (anonymous):

You can reduce this to 2(2x-9)=143

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So, without having to distribute the 2, since its already on the outside, you can just divide both sides by 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now you can add nine to both sides to get the x term by itself

OpenStudy (decentnabeel):

\[2\left(2x-5-4\right)=160-17\quad :\quad x=\frac{161}{4}\quad \left(\mathrm{Decimal:\quad }x=40.25\right)\]

OpenStudy (anonymous):

Yes, after you add the nine, divide both sides by 2 to get your final answer, 40.25

OpenStudy (anonymous):

ohhhh thankss guyss thanks for the explination

OpenStudy (anonymous):

turn up

OpenStudy (anonymous):

What is the solution to the following equation? 2(3x − 7) + 18 = 10 1 3 5 6

OpenStudy (decentnabeel):

\[2\left(3x-7\right)+18=10\quad :\quad x=1\] 2(3x − 7) + 18 = 10 6x-14+18=10 6x+4=10 subtract 4 on both sides 6x-4+4=10-4 6x=6 x=1

OpenStudy (decentnabeel):

are you understand @bby.nessa

OpenStudy (anonymous):

yea i understand thanks

OpenStudy (sohailiftikhar):

just take every terms without x to the right side of equation to get the value of x it's simple first try your own then ask for explanation that's the way of learning.